OpenStudy (anonymous):

graph the equation using the x. y-intercepts. x/5 - y/8= 1

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

x/5 - y/8 = 1 is the same as x/5 + y/(-8) = 1

jimthompson5910 (jim_thompson5910):

keep in mind that intercept form is x/a + y/b = 1 where 'a' is the x-intercept and 'b' is the y-intercept

OpenStudy (anonymous):

I am confused on how to solve it, if I can get an example possibly then I will be able to solve it

jimthompson5910 (jim_thompson5910):

say you had x/9 - y/13 = 1 that's the same as x/9 + y/(-13) it's in the form x/a + y/b = 1 where a = 9 and b = -13 so the x-intercept of x/9 - y/13 = 1 is 9, which is the point (9,0) and the y-intercept of x/9 - y/13 = 1 is -13, which is the point (0,-13)

OpenStudy (anonymous):

so then it would be (5,0) for x and (0,8) for y, right?

jimthompson5910 (jim_thompson5910):

close, b is not 8 though, it's -8

OpenStudy (anonymous):

right, thank you, Merry Christmas

jimthompson5910 (jim_thompson5910):

np